Find f'(x) of f(x)= sin{x/(x - sin[x/(x-sinx)])}


1
Expert's answer
2021-03-07T17:31:05-0500

Solution:

"f(x)=\sin\{{\dfrac{x}{x-\sin[\dfrac{x}{x-\sin x}]}}\}"

On differentiating both sides w.r.t. "x",

"f(x)=\cos\{{\dfrac{x}{x-\sin[\dfrac{x}{x-\sin x}]}}\}\times \{{\dfrac{x}{x-\sin[\dfrac{x}{x-\sin x}]}}\}'" [using chain rule]

"=\cos\{{\dfrac{x}{x-\sin[\dfrac{x}{x-\sin x}]}}\}\times \{{\dfrac{(x-\sin[\dfrac{x}{x-\sin x}])(x)'-x(x-\sin[\dfrac{x}{x-\sin x}])'}{(x-\sin[\dfrac{x}{x-\sin x}])^2}}\}" [Using quotient rule]

"=\cos\{{\dfrac{x}{x-\sin[\dfrac{x}{x-\sin x}]}}\}\times \{{\dfrac{(x-\sin[\dfrac{x}{x-\sin x}])-x(1-\cos[\dfrac{x}{x-\sin x}])[\dfrac{x}{x-\sin x}]'}{(x-\sin[\dfrac{x}{x-\sin x}])^2}}\}"

"=\cos\{{\dfrac{x}{x-\sin[\dfrac{x}{x-\sin x}]}}\}\times \{{\dfrac{(x-\sin[\dfrac{x}{x-\sin x}])-x(1-\cos[\dfrac{x}{x-\sin x}])[\dfrac{(x-\sin x)-x(1-\cos x)}{(x-\sin x)^2}]}{(x-\sin[\dfrac{x}{x-\sin x}])^2}}\}"

"=\cos\{{\dfrac{x}{x-\sin[\dfrac{x}{x-\sin x}]}}\}\times \{{\dfrac{(x-\sin[\dfrac{x}{x-\sin x}])-x(1-\cos[\dfrac{x}{x-\sin x}])[\dfrac{x\cos x-\sin x}{(x-\sin x)^2}]}{(x-\sin[\dfrac{x}{x-\sin x}])^2}}\}"

This is our required "f'(x)".